l8r:

i need help w my work Sample Question: Rewrite x1/2 in radical form. √x √x2 1/√x -√x

Aratox:

@hero

sllo:

Its √x

adrianluvvsyouu2:

If you mean x^(1/2) that would be the same as taking the square root

GoatyGoat:

@aratox

jhonyy9:

\[x^{1/2} = \sqrt{x}\] - in this way

l8r:

@goatygoat wrote:
@aratox
answer my dms plz

l8r:

@jhonyy9 wrote:
\[x^{1/2} = \sqrt{x}\] - in this way
ty

ykbatman:

tsk tsk dad

ykbatman:

i thought you were smart

grimreaper124:

lol

Gdub08:

\[\square \root\]x=x1/2